T1ASCII(1) T1ASCII(1) NAME t1ascii−converttype-1fontfrombinarytoASCII SYNOPSIS t1ascii[−llength][input[output]] DESCRIPTION t1asciiconvertsAdobeType1fontprogramsinPFB(binary)formattoPFA(hexadecimal)format.Ifthefileoutputisnotspecifiedoutputgoestothestandardoutput.Ifthefileinputisnot specifiedinputcomesfromthestandardinput. OPTIONS −−line−length=num,−lnum Setthemaximumlengthofencryptedlinesintheoutputtonum.(Thesearethelinesconsistingwhollyofhexadecimaldigits.)Thedefaultis64. SEEALSO t1binary(1),t1unmac(1),t1mac(1),t1asm(1),t1disasm(1) AdobeType1FontFormat AUTHORS LeeHetherington(ilh@lcs.mit.edu) EddieKohler(eddietwo@lcs.mit.edu) PortedtoMicrosoftC/C++CompilerandMS-DOSoperatingsystembyKai-UweHerbing(herb-ing@netmbx.netmbx.de). 5Mar1998 1